# GateLens: A Reasoning-Enhanced LLM Agent for Automotive Software Release Analytics

### v1

@@ 4.3. Robustness: Handling Out of Scope and Imprecise Queries (RQ2)
To assess the robustness of our approach in handling diverse user queries under real-world conditions, we conducted further experiments focusing on filtering out of scope queries as well as processing imprecise queries. For this purpose, the data analysis team at our industrial partner company manually selected 37 out of scope queries and 50 imprecise queries from the historical logs of the first-generation system, which are used to perform targeted evaluations.
@@
Out of scope queries are those that cannot be meaningfully answered using the available data. For example, a query like ”What is the most beautiful truck?” requires subjective judgment and cannot be resolved through database operations; it should be identified and filtered as out of scope. On the other hand, imprecise queries are those that can be answered using the database but contain ambiguous or inexact terms. For instance, a query such as ”Find some trucks for cases that are NOK” is considered imprecise because while it seeks truck names where test results are ”NOK” (failed), it uses ambiguous terminology - referring to ”trucks” instead of the actual database field ”name”, and mentions ”NOK” without specifying the ”test_result” field. Such imprecise queries require mapping informal language to precise database fields and conditions for proper execution.
4.3.1. Handling Out of Scope Queries
@@
We compared GateLens with other models; the results can be found in Table 3. The results demonstrate that GateLens with GPT-4o achieved the best performance, particularly in terms of precision, which is approximately 40% higher than other models, indicating GateLens’ ability to avoid generating incorrect results.
Table 3. Comparison of different models for out of scope queries.
@@
GoNoGo with GPT-4o 51.10% 89.19% 64.97%
The superior precision of GateLens with GPT-4o can be attributed to two key aspects of its design. First, its robust filtering mechanism ensures that out of scope queries are identified and excluded early in the processing pipeline, preventing irrelevant results. Second, the conversion of raw natural language queries into structured RA expressions enables the model to isolate and capture task-relevant components of a query. This structured approach considerably decreases erroneous outcomes and enhances the model’s ability to handle complex and diverse query formulations in real-world scenarios.
In contrast, the precision of GoNoGo was notably weaker. This can be attributed to the diversity of real-world queries and the variability in user narratives, which often contain a mixture of relevant and irrelevant parts. Such queries increase the degree of uncertainty, making it challenging for models that cannot effectively isolate task-relevant parts to process them appropriately. Furthermore, while the recall rates of all models were relatively high, indicating their ability to generate some results, this does not necessarily translate to accurate query processing.
A significant limitation was observed with GoNoGo, which relies on few-shot learning, as it occasionally generated code for out of scope queries. This behavior likely stems from the task-relevant examples in its prompt, which inadvertently bias the model toward attempting code generation even for out of scope queries. In summary, our approach demonstrates a strong ability to filter out of scope queries while maintaining a low error rate when processing in-scope queries.
4.3.2. Handling Imprecise Queries
@@
Table 4. Comparison of different models for imprecise queries.
@@
GoNoGo with GPT-4o 90% 36% 51.43%
The results of these experiments are presented in Table 4. As shown, GateLens with GPT-4 demonstrates the best overall performance. In terms of precision, all methods performed relatively well, indicating that when results are generated, they are likely to be correct. However, our method significantly outperformed the others in recall, highlighting its ability to handle a larger portion of the imprecise queries. As a result, GateLens with GPT-4 achieved a substantially higher F1 score compared to other methods, demonstrating that it not only processes most queries but also produces accurate results for them.
The observed performance gap between GateLens with GPT-4o and the other models can be attributed to their inherent limitations. Specifically, the Llama 3.1 70B model struggled to interpret user queries that deviated from the exact column header descriptions in the database schema. In such cases, Llama 3.1 70B often converted only the clearly defined parts of the query into RA, leading to incomplete query execution and reduced accuracy. On the other hand, GoNoGo exhibits low recall, as it is highly susceptible to confusion by ambiguous query elements. This causes GoNoGo to frequently generate incorrect code, which, when executed, fails to produce valid results, significantly lowering its recall rate.
@@ 4.4. Effectiveness of the RA module (RQ3)
Figure 3. Comparison of the original method and the method without the RA module across different datasets.
@@
4.5. Impact of Few-Shot Examples (RQ4)
To investigate the effect of including few-shot examples in prompts, we conducted experiments by varying the number of examples provided to both GateLens and GoNoGo. This experiment is performed on the first benchmark containing 50 designed queries, with results illustrated in Figure 4.
The results demonstrate that GateLens relies heavily on its RA translation process, achieving optimal performance even in a 0-shot setting without any examples. Interestingly, when a small number of examples are added, GateLens becomes slightly biased toward them, leading to a minor degradation in performance. However, it regains optimal performance with additional examples as the system learns to generalize better. In contrast, GoNoGo’s performance heavily depends on in-context examples, showing suboptimal results without sufficient few-shot examples.
The results further highlight the critical role of RA in enabling effective query handling. GateLens, through its RA-based translation process, achieves robust performance without requiring extensive in-context learning. GateLens’s independence from examples not only ensures consistent performance across diverse queries but also offers practical benefits by mitigating issues related to context length limitations and accordingly reducing computational and financial resource consumption, enhancing system scalability, and minimal maintenance needs for adaptation to new tasks. These advantages position GateLens as a robust and efficient solution for automating the analysis of tabular data in dynamic environments.
Figure 4. Performance of GateLens and GoNoGo with varying numbers of few-shot examples.
5. Industrial Deployment: Lesson Learned
Table 5. Comparison of GateLens (zero-shot) and GoNoGo (few-shot) performance across different roles on the second benchmark. For each role tested, GoNoGo was trained using examples from the other two roles only (leave-one-role-out approach).
Roles # Queries GateLens GoNoGo (Few-shot with All Roles) GoNoGo (Few-shot with Leave-One-Role-Out)
@@ 5. Industrial Deployment: Lesson Learned
The deployment we have done of GateLens in a partner automotive company highlights both the challenges and opportunities of integrating AI-assisted analytics into complex industrial workflows. The system was designed to streamline decision-making in automotive software release processes, and its deployment has provided valuable insights into improving decision-making in software release validation for large-scale environments.
Automotive software integration at the company typically occurs across three hierarchical stages: subsystem (control unit), system (multiple control units), and full vehicle levels. Each stage involves extensive testing, with results stored in a central database. Critical Go/No-Go decisions are made at these stages to determine whether a release meets quality thresholds. However, stakeholders from diverse backgrounds—including project managers, mechanical engineers, and software engineers—must query the raw data to evaluate product quality. Many lack expertise in data analytics, creating inefficiencies in decision-making.
Previously, these analytics were managed by a small team of 2–3 full-time analysts, who were often overwhelmed by the volume and diversity of requests. Scaling the team to meet the current demand would have required tripling its size. GateLens addresses this challenge by automating much of the workload, enabling decisions to be made more efficiently and accurately. Currently, GateLens is in an extended pilot phase, supporting a pool of 60-80 users. The analytics team has transitioned to a support role, helping stakeholders articulate their needs into clear, actionable prompts for the system.
User adoption of GateLens has progressed in phases:
  • Small-Scale Pilot: The initial deployment within the analytics team established baseline benchmarks.
@@
The diversity in stakeholders’ needs and backgrounds is an inevitable factor, leading to a wide range of requirements and queries when interacting with analytics systems. Our previous system, GoNoGo, has performed well when serving a specific group of stakeholders. However, as we opened the system to a broader audience, the variety of queries increased substantially. Systems that rely heavily on few-shot learning face significant challenges in these situations, as they are limited by the relatively small set of few-shot examples, making it difficult to handle a wider range of potentially unexpected inputs. Similarly, fine-tuned models often struggle to adapt to dynamic and diverse environments. Improving generalization is essential to ensure the scalability and reliability of analytics systems in these complex, domain-specific contexts. In our design and development work, we prioritized the system’s ability to generalize effectively and meet the needs of a broader and more diverse group of users.
@@
To explore the system’s generalizability, we categorized the roles within the company into three groups: mechanically-oriented, project-oriented, and software-oriented roles. The queries from these roles have distinct focuses. Mechanically-oriented roles typically focus on truck-specific data filtering. Project-oriented roles often combine meta-queries with conditional filters for release management and statistical analysis. Software roles emphasize truck software applications and user functions. We can see from Table 5, both GateLens (zero-shot) and GoNoGo (few-shot) exhibit differences in system performance across these groups, which is likely stemming from the complexity and variety of their typical queries. Nevertheless, the results demonstrate that GateLens is capable of supporting all groups to a high degree. To further evaluate GoNoGo’s dependency on few-shot examples, we conducted a leave-one-role-out experiment. In this approach, examples from a specific role are excluded in each iteration. For instance, ’without software’ indicates that all examples from the software-oriented role have been removed, while the total number of examples is maintained by substituting them with examples from other roles. This highlights the potential challenges with the robustness and generalizability of techniques that rely on few-shot examples. This is a crucial factor to consider in industries where diverse teams collaborate and a wide range of queries may arise.
@@
The impact of automated systems, such as GoNoGo and GateLens, on the release process has been substantial. Compared to the previous manual process, GateLens has reduced the time required for Go/No-Go analytics by more than 80%, significantly improving operational efficiency. Stakeholders can now focus on high-level decision-making, freed from the burden of data preparation and analysis.
@@
A key advantage of GateLens lies in its domain-specific design. Unlike general-purpose tools like TaskWeaver (Qiao et al., 2023), GateLens is tailored to automotive workflows, making it easier to understand, debug, and adapt to automotive common procedures. This focus on domain relevance ensures that the system aligns more closely with stakeholders’ needs while providing reliable and nuanced support.
@@ 6. Related Work
General-purpose LLMs are primarily designed for and trained on natural languages. Working with tabular data requires specialized adaptations to effectively handle its structured and heterogeneous nature (Fang et al., 2024; Wang et al., 2024b; van Breugel and van der Schaar, 2024). First, the structured tablular data is typically transformed into serialized text. The performance of the LLM may depend on this transformation (Min et al., 2024). Subsequently, the serialized text data is used as input to the LLM for various tasks, such as question-answering, summarization, or logical reasoning. Common solutions to improve the LLM performance on various tasks include prompting engineering, pre-training, fine-tuning, and Retrieval-Augmented Generation (RAG).
@@
Pre-training and fine-tuning (Zhang et al., 2023b; Parthasarathy et al., 2024; VM et al., 2024; Dong et al., 2022; Hegselmann et al., [n. d.]) often face scalability concerns. Athough resource-efficient training techniques have been proposed to mitigate the substantial computational demands of LLMs (Han et al., 2024; Lin et al., 2024), in safety-critical applications with evolving data and requirements, training LLMs presents significant challenges due to the constant need for rigorous validation and verification. This ongoing necessity substantially increases resource demands for development and maintenance, potentially exceeding the capacities of many companies. Techniques such as RAG have been employed to dynamically integrate external knowledge bases during inference, reducing the need for frequent model updates (Zhao et al., 2024; Gao et al., 2023). However, such methods can pose challenges in safety-critical industrial settings as well, since both retrieval modules and model components must undergo synchronized updates to maintain relevance, reliability, and compliance with validation and verification requirements.
@@
Prompt engineering techniques are among the most resource-efficient methods for improving LLM output (Sahoo et al., 2024; Jin and Lu, 2023). From a user standpoint, when the input is natural language, prompting techniques can be broadly categorized based on the type of language generated by the LLM. These include outputs in natural language, structured languages (Li et al., 2023), or symbolic languages. When the generated language is natural language, LLMs often fail to consistently follow instructions, particularly when the instructions are complex or require precise, step-by-step execution (Pham et al., 2024). This inconsistency arises because natural language, while flexible and expressive, can be ambiguous and prone to misinterpretation by LLMs. Structured languages include general-purpose languages (e.g. Python) (Ye et al., 2024), query languages (e.g. SQL) (Li et al., [n. d.]; Dong et al., 2023; Mouravieff et al., 2024), configuration formats (e.g. YAML or JSON), or other Domain-Specific Languages (DSLs). These lagnguages are subsequently interpreted and/or executed by either external tools, the same LLM, or another LLM agent. This approach offers significant advantages, as it enables precise execution of tasks. Another popular type of output is symbolic languages. Literature shows that symbolic representations provide a more rigorous framework for articulating premises and intent, which can enhance reasoning capabilities (Pan et al., 2023).
In this paper, we uniquely integrate structured and symbolic languages by converting natural language input into RA expressions, a formalism specifically designed for relational modeling and particularly suited for table analysis, before translating these expressions into Python code for execution. This approach contrasts with prior work that often relies on specialized DSLs or fine-tuning, which can limit flexibility and scalability. By leveraging RA as an intermediate symbolic representation, we achieve precise reasoning and query optimization, while Python’s general-purpose capabilities ensure broad applicability and efficient execution across diverse scenarios. Unlike methods that depend heavily on few-shot examples or fine-tuning, GateLens combines interpretability, adaptability, and computational efficiency, making it particularly suitable for dynamic, safety-critical table analysis applications in industrial environments. Notably, RAG systems focus on factuality by retrieving concrete examples without dedicated reasoning modules; nonetheless, they can be integrated with GateLens if needed. It is worth noting that RAG requires constant updates and rigorous validation for safety-critical applications. This ongoing process demands significant effort, specialized expertise, and substantial resources, which can pose a considerable challenge for some organizations.
7. Discussion and Future Work
This study introduced GateLens, a reasoning-enhanced LLM agent specifically designed for automotive software release validation. GateLens demonstrates significant advancements by reducing analysis time by over 80% while maintaining high accuracy in test result interpretation, impact assessment, and release candidate evaluation. The inclusion of a Relational Algebra reasoning module is a critical factor in improving robustness and scalability, as evidenced by superior F1 scores in both benchmarking and industrial evaluations. Additionally, GateLens successfully handles complex and ambiguous queries, providing practical and reliable support for faster and more informed decision-making in safety-critical software release processes. A phased deployment program is ongoing and shows that multiple stakeholder groups can be supported but also that the roles change and analytical needs evolve which will require continued refinement and discussions.
The validity of our findings is subject to several potential threats. First, the generalizability of GateLens beyond the automotive software release domain is limited, as the system relies on domain-specific relational modeling and datasets, necessitating further validation in other safety-critical industries. Second, the benchmarks and query scenarios used for evaluation, derived from historical data and real-world queries, may not fully capture the diversity and complexity of potential use cases, which could impact the robustness of the system in broader deployments. Finally, the system’s performance is influenced by the specific configurations of the LLMs used, such as GPT-4o and Llama 3.1 70B, requiring the capability of interpreting and generating RA expressions. Future work will address these limitations by expanding domain applicability, diversifying evaluation datasets, and testing alternative LLM configurations to enhance reliability and scalability.
8. Conclusions
In automotive manufacturing, efficient software release and status monitoring depend critically on analyzing test results logged in a database. However, the volume and diversity of required queries from verification engineers and stakeholders make manual analysis a significant challenge. GateLens addresses these challenges by leveraging an LLM-based approach that translates natural language queries into structured relational algebra expressions within a defined relational model, in which the agent uses these expressions to produce the desired analytical code. Besides, GateLens has the ability to filter out-of-scope queries based on the database schema, improving the robustness and reliability of the framework.

### v2

@@ IV-C Robustness: Handling Out of Scope and Imprecise Queries (RQ2)
To assess the robustness of our approach in handling diverse user queries under real-world conditions, we conducted further experiments focusing on filtering out-of-scope queries as well as processing imprecise queries. For this purpose, the data analysis team at our industrial partner company manually selected 37 out-of-scope queries and 50 imprecise queries from the historical logs of the first-generation system, which are used to perform targeted evaluations.
@@
Out-of-scope queries are those that cannot be meaningfully answered using the available data. For example, a query like ”What is the most beautiful truck?” requires subjective judgment and cannot be resolved through database operations; it should be identified and filtered as out of scope. On the other hand, imprecise queries are those that can be answered using the database but contain ambiguous or inexact terms. For instance, a query such as ”Find some trucks for cases that are NOK” is considered imprecise because while it seeks truck names where test results are ”NOK” (failed), it uses ambiguous terminology - referring to ”trucks” instead of the actual database field ”name”, and mentions ”NOK” without specifying the ”test_result” field. Such imprecise queries require mapping informal language to precise database fields and conditions for proper execution.
IV-C1 Handling Out of Scope Queries
@@
We compared GateLens with other models; the results can be found in Table III. The results demonstrate that GateLens with GPT-4o achieved the best performance, particularly in terms of precision, which is approximately 40% higher than other models, indicating GateLens’ ability to avoid generating incorrect results.
The superior precision of GateLens with GPT-4o can be attributed to two key aspects of its design. First, its robust filtering mechanism ensures that out-of-scope queries are identified and excluded early in the processing pipeline, preventing irrelevant results. Second, the conversion of raw natural language queries into structured RA expressions enables the model to isolate and capture task-relevant components of a query. This structured approach considerably decreases erroneous outcomes and enhances the model’s ability to handle complex and diverse query formulations in real-world scenarios.
TABLE III: Model comparison for out-of-scope queries.
@@
CoT+SC with GPT-4o 51.10% 89.19% 64.97%
CoT+SC showed significantly lower precision due to the variability of real-world queries and the inconsistency of user narratives, which often contain a mix of relevant and irrelevant content. This variability increases uncertainty and poses challenges for models that struggle to identify task-relevant information. Although all models demonstrated high recall, this did not translate into accurate processing.
IV-C2 Handling Imprecise Queries
@@
TABLE IV: Model comparison for imprecise queries.
@@
CoT+SC with GPT-4o 90% 36% 51.43%
The results of these experiments are presented in Table IV. As shown, GateLens with GPT-4o demonstrates the best overall performance. In terms of precision, all methods performed relatively well, indicating that when results are generated, they are likely to be correct. However, our method significantly outperformed the others in recall, highlighting its ability to handle a larger portion of the imprecise queries. As a result, GateLens with GPT-4 achieved a substantially higher F1 score compared to other methods, demonstrating that it not only processes most queries but also produces accurate results for them.
The observed performance gap between GateLens with GPT-4o and the other models can be attributed to their inherent limitations. Specifically, the Llama 3.1 70B model struggled to interpret user queries that deviated from the exact column header descriptions in the database schema. In such cases, Llama 3.1 70B often converted only the clearly defined parts of the query into RA, leading to incomplete execution and reduced accuracy. On the other hand, CoT+SC exhibits low recall, as it is highly susceptible to confusion by ambiguous query elements. This causes CoT+SC to frequently generate incorrect code that fails execution, significantly lowering its recall rate.
@@ IV-D Effectiveness of the RA module (RQ3)
Figure 3: Comparison of the original method and the method without the RA module across different datasets.
@@
IV-E Impact of Few-Shot Examples (RQ4)
To investigate the effect of including few-shot examples in prompts, we conducted experiments by varying the number of examples provided to both GateLens and CoT+SC. This experiment is performed on the first benchmark containing 50 designed queries, with results illustrated in Figure 4.
The results demonstrate that GateLens relies heavily on its RA translation process, achieving optimal performance even in a 0-shot setting without any examples. Interestingly, when a small number of examples are added, GateLens becomes slightly biased toward them, leading to a minor degradation in performance. However, it regains optimal performance with additional examples as the system learns to generalize better. In contrast, CoT+SC’s performance heavily depends on in-context examples, showing suboptimal results without sufficient few-shot examples. While CoT+SC’s performance improves with more examples, this approach is fundamentally inefficient and impractical. Increasing example count expands input context size, leading to higher latency due to the quadratic complexity of Transformer-based models, particularly problematic for real-time applications. This also escalates operational costs through increased token usage (higher cloud API fees) and computational demands. Moreover, the larger context risks exceeding the maximum length, which can lead to truncation or lost information, compromising the model’s response quality. In long contexts, critical content may be ignored, resulting in a phenomenon known as “lost in the middle,” which adversely affects overall performance (Liu et al., 2023).
The results further highlight the critical role of RA in enabling effective query handling. GateLens, through its RA-based translation process, achieves robust performance without requiring extensive in-context learning. GateLens’s independence from examples not only ensures consistent performance across diverse queries but also offers practical benefits by mitigating issues related to context length limitations and accordingly reducing computational and financial resource consumption, enhancing system scalability, and minimizing maintenance needs for adaptation to new tasks. These advantages position GateLens as a robust and efficient solution for automating the analysis of tabular data in dynamic environments.
Figure 4: Comparison of GateLens against CoT+SC across different numbers of few-shot examples.
V Industrial Deployment: Lessons Learned
TABLE V: GateLens (zero-shot) vs CoT+SC (few-shot) performance across different roles on the second benchmark. For each role tested, CoT+SC was trained using examples from the other two roles only (leave-one-role-out approach).
Roles # Queries GateLens CoT+SC (Few-shot with All Roles) CoT+SC (Few-shot with Leave-One-Role-Out)
@@ V Industrial Deployment: Lessons Learned
The deployment of GateLens at a partner automotive company has provided valuable insights into integrating AI-assisted analytics into complex industrial workflows, specifically for streamlining decision-making in automotive software release validation.
Automotive software integration at the company typically occurs across three hierarchical stages: subsystem (control unit), system (multiple control units), and full vehicle levels. Each stage involves extensive testing, with results stored in a central database. Critical Go/No-Go decisions are made at these stages to determine whether a release meets quality thresholds. However, stakeholders from diverse backgrounds—including project managers, mechanical engineers, and software engineers—must query the raw data to evaluate product quality. Many lack expertise in data analytics, creating bottlenecks and delays in the decision-making process.
Previously, these analytics were managed by a small team of 2–3 full-time analysts, who were often overwhelmed by the volume and diversity of requests. Scaling the team to meet the current demand would have required tripling its size. GateLens addresses this challenge by automating much of the workload, enabling more efficient decision-making. Currently, GateLens is in an extended pilot phase, supporting a pool of 60-80 users. The analytics team has transitioned to a support role, helping stakeholders articulate their needs into clear, actionable prompts for the system. User adoption of GateLens has progressed in phases:
  • Small-Scale Pilot: The initial deployment within the analytics team established benchmarks.
@@
The diversity in stakeholders’ needs and backgrounds is an inevitable factor, leading to a wide range of requirements and queries when interacting with analytics systems. Our preliminary agentic system based on CoT and SC has performed well when serving a specific group of stakeholders. However, as we opened the system to a broader audience, the variety of queries increased substantially. Systems that rely heavily on few-shot learning face significant challenges in these situations, as they are limited by the relatively small set of few-shot examples, making it difficult to handle a wider range of potentially unexpected inputs. Similarly, fine-tuned models often struggle to adapt to dynamic and diverse environments. Improving generalization is essential to ensure the scalability and reliability of analytics systems in these complex, domain-specific contexts. In our design and development work, we prioritized the system’s ability to generalize effectively and meet the needs of a broader and more diverse group of users.
@@
To explore the system’s generalizability, we categorized the roles within the company into three groups: mechanically-oriented, project-oriented, and software-oriented roles. Mechanically-oriented roles typically focus on truck-specific data filtering. Project-oriented roles often combine meta-queries with conditional filters for release management and statistical analysis. Software roles emphasize truck software applications and user functions. We can see from Table V, both GateLens (zero-shot) and CoT+SC (few-shot) exhibit differences in system performance across these groups, which is likely stemming from the complexity and variety of their typical queries. Nevertheless, the results demonstrate that GateLens is capable of supporting all groups to a high degree. To further evaluate CoT+SC’s dependency on few-shot examples, we conducted a leave-one-role-out experiment. In this approach, examples from a specific role are excluded in each iteration. For instance, ’without software’ indicates that all examples from the software-oriented role have been removed, while the total number of examples is maintained by substituting them with examples from other roles. This highlights the potential challenges with the robustness and generalizability of techniques that rely on few-shot examples. This is a crucial factor to consider in industries where diverse teams collaborate and a wide range of queries may arise.
@@
The impact of automated systems, such as GateLens, on the release process has been substantial. Compared to the previous manual process, GateLens has reduced the time required for Go/No-Go analytics by more than 80%, significantly improving operational efficiency. Stakeholders can now focus on high-level decision-making, freed from the burden of data preparation and analysis.
@@
A key advantage of GateLens lies in its domain-specific design. Unlike general-purpose tools like TaskWeaver (Qiao et al., 2023) or AutoGen (Wu et al., 2023), GateLens is tailored to automotive workflows, making it easier to understand, debug, and adapt to automotive common procedures. This focus on domain relevance ensures that the system aligns more closely with stakeholders’ needs while providing reliable and nuanced support.
@@ VI Related Work
General-purpose LLMs are primarily designed for and trained on natural languages. Working with tabular data requires specialized adaptations to effectively handle its structured and heterogeneous nature (Fang et al., 2024; Wang et al., 2024b; van Breugel and van der Schaar, 2024). First, the structured tabular data is typically transformed into serialized text. The performance of the LLM may depend on this transformation (Min et al., 2024). Subsequently, the serialized text data is used as input to the LLM for various tasks, such as question-answering, summarization, or logical reasoning. Common approaches to improve LLM performance include prompt engineering, pre-training, fine-tuning, and Retrieval-Augmented Generation (RAG).
@@
Pre-training and fine-tuning (Zhang et al., 2023b; Parthasarathy et al., 2024; VM et al., 2024; Dong et al., 2022; Hegselmann et al., ) often face scalability concerns. Although resource-efficient training techniques have been proposed to mitigate the substantial computational demands of LLMs (Han et al., 2024; Lin et al., 2024), in safety-critical applications with evolving data and requirements, training LLMs presents significant challenges due to the constant need for rigorous validation and verification. This ongoing necessity substantially increases resource demands for development and maintenance, potentially exceeding the capacities of many companies. Techniques such as RAG have been employed to dynamically integrate external knowledge bases during inference, reducing the need for frequent model updates (Zhao et al., 2024; Gao et al., 2023). However, such methods can pose challenges in safety-critical industrial settings as well, since both retrieval modules and model components must undergo synchronized updates to maintain relevance, reliability, and compliance with validation and verification requirements. Costs would also be especially high with fine-tuning since re-tuning would be needed when new and improved base LLMs are released and should be incorporated.
@@
Prompt engineering techniques are among the most resource-efficient methods for improving LLM output (Sahoo et al., 2024; Jin and Lu, 2023). From a user standpoint, when the input is natural language, prompting techniques can be broadly categorized based on the type of language generated by the LLM. These include outputs in natural language, structured languages (Li et al., 2023), or symbolic languages. When the generated language is natural language, LLMs often fail to consistently follow instructions, particularly when the instructions are complex or require precise, step-by-step execution (Pham et al., 2024). This inconsistency arises because natural language, while flexible and expressive, can be ambiguous and prone to misinterpretation by LLMs. Structured languages include general-purpose languages (e.g. Python) (Ye et al., 2024), query languages (e.g. SQL) (Li et al., ; Dong et al., 2023; Mouravieff et al., 2024), configuration formats (e.g. YAML or JSON), or other Domain-Specific Languages (DSLs) (Glenn et al., 2024; Dai et al., 2024). These languages are subsequently interpreted and/or executed by either external tools, the same LLM, or another LLM agent. This approach offers significant advantages, as it enables precise execution of tasks. Another popular type of output is symbolic languages. Literature shows that symbolic representations provide a more rigorous framework for articulating premises and intent, which can enhance reasoning capabilities (Pan et al., 2023).
In this paper, we introduce a novel prompt-only (training-free) approach that bridges natural language and executable code through RA, a symbolic formalism designed for relational modeling and ideally suited for analyzing tabular data. Unlike prior work that often relies on complex multi-agent planning, our approach leverages RA as a lightweight intermediate representation to enable precise query normalization, disambiguation of natural language input, and efficient code generation. RA acts as an abstraction layer that can target multiple execution backends (e.g., Python, SQL), providing adaptability across systems. In GateLens, we generate Python code to support practical industrial deployment and high-performance execution. GateLens is training-free, feed-forward (single-pass, without looping or multi-agent orchestration), and thus easier to verify, trace, maintain, and trust – qualities critical for safety-critical industrial applications.
VII Discussion and Conclusions
This study introduced GateLens, a reasoning-enhanced LLM agent designed for automotive software release validation. By introducing RA as an intermediate representation before code generation, Gatelens addresses the ”Unfaithful Chain-of-Thought (CoT) reasoning” problem in code generation, where reasoning steps in CoT explanations do not accurately reflect the model’s actual thought process (Turpin et al., 2023). Specifically, GateLens divides the analysis into two steps: (1) natural language queries are first translated into RA expressions, and (2) these RA expressions are then converted into executable code. We use Python as our target language in step (2) due to its widespread use in our partner company and the model’s strong performance in Python, which benefits from more extensive training data. However, this step is also compatible with RA-to-SQL generation, enabling flexibility across backends. The inclusion of the RA reasoning module is a critical factor in improving robustness and scalability, as evidenced by superior F1 scores in both benchmarking and industrial evaluations.